OpenStudy (anonymous):

Solve cos x + 2√ = -cos x for x over the interval [0,2Π)

OpenStudy (freckles):

equation seems weird square root of an equal sign?

OpenStudy (jdoe0001):

\(\large cos(x)+\sqrt{2}=-cos(x)?\)

OpenStudy (xapproachesinfinity):

seems good with jdoe correction

OpenStudy (xapproachesinfinity):

add cosx to both sides then subtract root2 from both sides

OpenStudy (xapproachesinfinity):

getting \[2\cos x=\sqrt{2}\]

OpenStudy (xapproachesinfinity):

see what can you do now